Introduction

The best choice problem for random number of

objects with a refusal probability

M Horiguchi; Kanagawa Univeresity, Japan M Yasuda; Chiba University, Japan

July 2009, Cornell

The best choice problem for random number of objects with a refusal probability

Abstract:

From the motivation to make a material for teaching Math & Stat, some of usful software are introduced. We mainly tried as follow;

(1) e-math (2) ketpic (3) wintpic.

We hope that student understand, corporating mathematical expression or equation, visual computer display and explicit graph based on Compter Algebra Sytem.

The best choice problem for random number of objects with a refusal probability

Keyword:

Computer Algebra Systems, ⇒ Mathematica, Maple, SciLab. Macro packages:

▶ e-math ⇒ style file of latex.

▶ ketpic ⇒ tpic form combined with CAS.

▶ wintpic ⇒ tpic form with (half)WYSWYG. Show samples of figures, text material in math etc... : Conference and Journal: ”Future Generation Computer Systems”, Copyright 2008 Elsevier B.V. International Workshop on Computer Graphics and Geometric Modeling

- Krakow (Poland), June 23-25 2008. Technical Session on Computer Graphics and Geometric Modeling - TSCG, Kuala Lumpur (Malaysia), August 27-30 2007.

Computer Algebra Systems, ⇒ Mathematica, Maple, SciLab. Macro packages:

▶ e-math ⇒ style file of latex.

▶ ketpic ⇒ tpic form combined with CAS.

▶ wintpic ⇒ tpic form with (half)WYSWYG. Show samples of figures, text material in math etc... : Conference and Journal: ”Future Generation Computer Systems”, Copyright 2008 Elsevier B.V. International Workshop on Computer Graphics and Geometric Modeling

- Krakow (Poland), June 23-25 2008. Technical Session on Computer Graphics and Geometric Modeling - TSCG, Kuala Lumpur (Malaysia), August 27-30 2007.

Computer Algebra Systems, ⇒ Mathematica, Maple, SciLab. Macro packages:

▶ e-math ⇒ style file of latex.

▶ ketpic ⇒ tpic form combined with CAS.

▶ wintpic ⇒ tpic form with (half)WYSWYG. Show samples of figures, text material in math etc... : Conference and Journal: ”Future Generation Computer Systems”, Copyright 2008 Elsevier B.V. International Workshop on Computer Graphics and Geometric Modeling

- Krakow (Poland), June 23-25 2008. Technical Session on Computer Graphics and Geometric Modeling - TSCG, Kuala Lumpur (Malaysia), August 27-30 2007.